Grade 12Physical Chemistry

Calculate the time in which the activity of an element is reduced to 90% of its original value. The half-life period of the element is 1.4*10^10 years . Answer :- 2.13*10^9 years

Harshdeep Kaur
11 Years agoGrade 12

1 Answer

Sunil Kumar FP
11 Years ago
let the initial activity of the element be 100
final activity of the element=10
we have
k=2.303./T *log (initial concentration/final concentration)
k=decay constant=.693/t1/2)
.693/1.4*10^10 =2.303/T log(100/90)
T=2.303*1.4*10^10*.045/.693
=2.1*10^9